Four temperature readings in °C for Mrs. Baskin are 37.6, 35.5, 38.1, and 37.6. Find her average
(or mean) °C temperature and convert it to °F. Average: __________ °C ? __________ °F
Answer #1

ANSWER

     Average or mean temperature of Mr. Baskin.

       * Temperature reading of Mr.Baskin = 37.6, 35.5 , 38.1, 37.6

              Formula = Sum of all temperature readings / Total number of temperature readings.

                                = 37.6 + 35.5 + 38.1 + 37.6 /4

                                = 148.8 / 4

                                = 37.2 Degree celcious.

* Convert above reading to fahrenheit

                 Formula

                       Temperature in fahrenheit = Temperature in celcious * 9/5 +32

                                                                 = 37.2 * 9/5 +32

                                                                 = 66.96 +32

                                                                = 98.96 Fahreheit